# Odd-parity Magnetism Driven by Antiferromagnetic Exchange

### v2

@@ Appendix E List of materials
[…]
@@
$C_{2h}$ $B_{g}$ $\begin{matrix}c_{1}k_{x}k_{z}\\ +c_{2}k_{y}k_{z}\end{matrix}$ $\begin{matrix}11(C^{*}_{1},D^{*}_{1},E^{*}_{1},Z^{*}_{1}),13(A^{*}_{1},B^{*}_{1},D^{*}_{1},E^{*}_{1}),14(A^{*}_{1},B^{*}_{1},C^{*}_{1},Z^{*}_{1}),15(A^{*}_{1},M^{*}_{1}),63(R_{1}),\\ 64(S_{1}),67(R_{1},S_{1}),68(R_{1},S_{1}),72(S^{*}_{1},R^{*}_{1}),73(T_{1},S^{*}_{1},R^{*}_{1}),74(T_{1}),85(R_{1},X_{1}),\\ 86(R_{1},X_{1}),88(X_{1}),140(N^{*}_{1}),142(N^{*}_{1}),163(L^{*}_{1}),165(L^{*}_{1}),167(L^{*}_{1}),176(L_{1}),206(N_{1})\end{matrix}$
$C_{4h}$ $B_{g}$ $\begin{matrix}c_{1}(k_{x}^{2}-k_{y}^{2})\\ +c_{2}k_{x}k_{y}\end{matrix}$ $84(A_{1,2},Z_{1,2}),85(A_{1,2},M_{1,2}),86(M_{1,2},Z_{1,2}),88(M_{1,2})$
@@
$D_{2h}$ $\begin{matrix}B_{1g}\\ B_{2g}\\ B_{3g}\end{matrix}$ $\begin{matrix}k_{x}k_{y}\\ k_{z}k_{x}\\ k_{y}k_{z}\end{matrix}$ $\begin{matrix}48(S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),49(R_{1,2},T_{1,2},U_{1,2},Z_{1,2}),\\ 50(R_{1,2},S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),51(R_{1,2},S_{1,2},U_{1,2},X_{1,2}),52(R_{1,2},U_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),\\ 53(S_{1,2},T_{1,2},X_{1,2},Z_{1,2}),54(S_{1,2},T_{1,2},X_{1,2},Z_{1,2}),55(T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),\\ 56(S_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),57(S_{1,2},U_{1,2},Y_{1,2},Z_{1,2}),58(R_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),\\ 59(R_{1,2},S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),60(X_{1,2},Y_{1,2},Z_{1,2}),61(X_{1,2},Y_{1,2},Z_{1,2}),\\ 62(T_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),63(T_{1,2},Z_{1,2}),64(T_{1,2},Z_{1,2}),66(T_{1,2},Z_{1,2}),68(T_{1,2},Z_{1,2}),\\ 70(T_{1,2},Y_{1,2},Z_{1,2}),124(R_{1,2}),125(R_{1,2},X_{1,2}),126(R_{1,2},X_{1,2}),127(R_{1,2},X_{1,2}),128(X_{1,2}),\\ 129(R_{1,2},X_{1,2}),130(X_{1,2}),132(R_{1,2}),133(R_{1,2},X_{1,2}),134(R_{1,2},X_{1,2}),135(R_{1,2},X_{1,2}),\\ 136(X_{1,2}),137(R_{1,2},X_{1,2}),138(X_{1,2}),141(X_{1,2}),142(X_{1,2}),\\ 192(L_{1,2}),193(L_{1,2}),194(L_{1,2}),201(M_{1,2},X_{1,2}),203(X_{1,2}),205(X_{1,2}),230(N_{1,2})\end{matrix}$
@@
$D_{4h}$ $B_{1g}$ $k_{x}^{2}-k_{y}^{2}$ $\begin{matrix}126(A_{1,2,3,4}),131(A_{1,2,3,4},Z_{1,2,3,4}),133(Z_{1,2,3,4}),135(Z_{1,2,3,4}),\\ 137(Z_{1,2,3,4}),141(M_{1,2,3,4}),142(M_{1,2,3,4}),223(X^{*}_{1,2,3,4})\end{matrix}$
$D_{4h}$ $B_{2g}$ $k_{x}k_{y}$ $\begin{matrix}125(A_{1,2,3,4},M_{1,2,3,4}),126(M_{1,2,3,4}),129(A_{1,2,3,4},M_{1,2,3,4}),130(M_{1,2,3,4}),\\ 132(A_{1,2,3,4},Z_{1,2,3,4}),133(M_{1,2,3,4}),134(M_{1,2,3,4},Z_{1,2,3,4}),136(A_{1,2,3,4},Z_{1,2,3,4}),\\ 137(M_{1,2,3,4}),138(M_{1,2,3,4},Z_{1,2,3,4}),222(M_{1,2,3,4}),\\ 224(M_{1,2,3,4},X^{*}_{1,2,3,4}),227(X^{*}_{1,2,3,4}),228(X^{*}_{1,2,3,4})\end{matrix}$
@@
$D_{2h}$ $A_{u}$ $xyz$ $\begin{matrix}48(S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),49(R_{1,2},T_{1,2},U_{1,2},Z_{1,2}),\\ 50(R_{1,2},S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),52(U_{1,2},X_{1,2},Z_{1,2}),53(S_{1,2},X_{1,2}),54(T_{1,2},Z_{1,2}),\\ 56(Z_{1,2}),58(Z_{1,2}),60(Y_{1,2}),66(T_{1,2},Z_{1,2}),68(T_{1,2},Z_{1,2}),70(T_{1,2},Y_{1,2},Z_{1,2}),124(R_{1,2}),\\ 125(R_{1,2},X_{1,2}),126(R_{1,2},X_{1,2}),132(R_{1,2}),133(R_{1,2},X_{1,2}),134(R_{1,2},X_{1,2}),\\ 141(X^{*}_{1,2}),142(X^{*}_{1,2}),192(L_{1,2}),201(M_{1,2},X_{1,2}),203(X_{1,2}),230(N^{*}_{1,2})\end{matrix}$
$D_{4h}$ $B_{1u}$ $xyz$ $\begin{matrix}125(A_{3,4},M_{3,4}),126(M_{3,4}),132(A_{1,2},Z_{1,2}),133(M_{1,2}),134(M_{1,2},Z_{1,2}),\\ 136(Z_{1,2}),138(Z_{1,2}),222(M_{3,4}),224(M_{1,2},X^{*}_{1,2}),227(X^{*}_{1,2}),228(X^{*}_{1,2})\end{matrix}$
@@ Appendix F Electronic structure of FeSe
Density-functional theory calculations for bulk FeSe with the $(\pi,\pi,0)$ coplanar magnetic order are carried out using the Full-potential Linearized Augmented Plane Wave (FLAPW) method Weinert et al. (2009), with the scalar-relativistic approximation Koelling and Harmon (1977); MacDonald et al. (1980). The Perdew-Burke-Ernzerhof form of exchange correlation functional Perdew et al. (1996) is used, along with wave function and potential energy cutoffs of 16 and 200 Ry, respectively. A muffin-tin sphere radius of 1.1 Å is applied to both Fe and Se atoms. The superlattice setup is $\sqrt{2}\times\sqrt{2}\times 1$, as indicated by the dotted line in the left panel of Figure 2 of the main text, with experimental lattice parameters ($a=3.765$; $c=5.518$; $z_{\textrm{Se}}=0.26$) Villars (1997). The Brillouin zone is sampled using (15,15,10) and (100,100,50) k-point meshes for self-consistent field calculation and Fermi surface calculation, respectively. The resulting band structure agrees with a previous calculation employing the generalized Bloch theorem for spin spirals Shishidou et al. (2018).
@@ Appendix G Electronic structure of CeNiAsO
For the construction of the minimal low-energy models relevant for CeNiAsO, we perform ab initio calculations within the full-potential local-orbital (FPLO) code Koepernik and Eschrig (1999) (version 22.00-62) in the non-relativistic setting and the LDSA approximation. We use $a=b=4.014537\mathrm{\AA}$, $c=7.90723\mathrm{\AA}$ in space group 129, Ce on 2c with internal coordinates $(1/4,1/4,0.154739)$, Ni on 2b with $(1/4,1/4,1/2)$, As on 2c with $(1/4,1/4,-0.351368)$ and O on 2a with $(1/4,3/4,0)$. The resulting band structure is shown in Fig. 5 which exhibits two pairs of not hybridized bands as checked by calculating the irreps of the corresponding bands.

### v3

@@ Appendix E List of materials
Table 2: Little co-group, even-parity order, space group and wavevectors. Labels denote orders not along the conventional axis. For instance, $1C_{1}$ hosts nematicity $c_{1}k_{x}k_{y}+c_{2}k_{z}k_{y}$.
@@
$C_{2h}$ $B_{g}$ $\begin{matrix}c_{1}k_{x}k_{z}\\ +c_{2}k_{y}k_{z}\end{matrix}$ $\begin{matrix}11(Z^{*}_{1},C^{*}_{1},D^{*}_{1},E^{*}_{1}),13(B^{*}_{1},D^{*}_{1},A^{*}_{1},E^{*}_{1}),14(Z^{*}_{1},B^{*}_{1},C^{*}_{1},A^{*}_{1}),15(A^{*}_{1},M^{*}_{1}),63(R_{1}),\\ 64(S_{1}),67(S_{1},R_{1}),68(S_{1},R_{1}),72(S^{*}_{1},R^{*}_{1}),73(S^{*}_{1},R^{*}_{1},T_{1}),74(T_{1}),85(X_{1},R_{1}),\\ 86(X_{1},R_{1}),88(X_{1}),140(N^{*}_{1}),142(N^{*}_{1}),163(L^{*}_{1}),165(L^{*}_{1}),167(L^{*}_{1}),176(L_{1}),206(N_{1})\end{matrix}$
$C_{4h}$ $B_{g}$ $\begin{matrix}c_{1}(k_{x}^{2}-k_{y}^{2})\\ +c_{2}k_{x}k_{y}\end{matrix}$ $84(Z_{1,2},A_{1,2}),85(M_{1,2},A_{1,2}),86(Z_{1,2},M_{1,2}),88(M_{1,2})$
@@
$D_{2h}$ $\begin{matrix}B_{1g}\\ B_{2g}\\ B_{3g}\end{matrix}$ $\begin{matrix}k_{x}k_{y}\\ k_{z}k_{x}\\ k_{y}k_{z}\end{matrix}$ $\begin{matrix}48(X_{1,2},Z_{1,2},U_{1,2},Y_{1,2},S_{1,2},T_{1,2}),49(Z_{1,2},U_{1,2},T_{1,2},R_{1,2}),\\ 50(X_{1,2},U_{1,2},Y_{1,2},S_{1,2},T_{1,2},R_{1,2}),51(X_{1,2},U_{1,2},S_{1,2},R_{1,2}),52(X_{1,2},Z_{1,2},U_{1,2},Y_{1,2},R_{1,2}),\\ 53(X_{1,2},Z_{1,2},S_{1,2},T_{1,2}),54(X_{1,2},Z_{1,2},S_{1,2},T_{1,2}),55(X_{1,2},U_{1,2},Y_{1,2},T_{1,2}),\\ 56(X_{1,2},Z_{1,2},Y_{1,2},S_{1,2}),57(Z_{1,2},U_{1,2},Y_{1,2},S_{1,2}),58(X_{1,2},Z_{1,2},Y_{1,2},R_{1,2}),\\ 59(X_{1,2},U_{1,2},Y_{1,2},S_{1,2},T_{1,2},R_{1,2}),60(X_{1,2},Z_{1,2},Y_{1,2}),61(X_{1,2},Z_{1,2},Y_{1,2}),\\ 62(X_{1,2},Z_{1,2},Y_{1,2},T_{1,2}),63(T_{1,2},Z_{1,2}),64(T_{1,2},Z_{1,2}),66(T_{1,2},Z_{1,2}),68(T_{1,2},Z_{1,2}),\\ 70(T_{1,2},Z_{1,2},Y_{1,2}),124(R_{1,2}),125(R_{1,2},X_{1,2}),126(R_{1,2},X_{1,2}),127(R_{1,2},X_{1,2}),128(X_{1,2}),\\ 129(R_{1,2},X_{1,2}),130(X_{1,2}),132(R_{1,2}),133(R_{1,2},X_{1,2}),134(R_{1,2},X_{1,2}),135(R_{1,2},X_{1,2}),\\ 136(X_{1,2}),137(R_{1,2},X_{1,2}),138(X_{1,2}),141(X^{*}_{1,2}),142(X^{*}_{1,2}),\\ 192(L^{*}_{1,2}),193(L^{*}_{1,2}),194(L^{*}_{1,2}),201(M_{1,2},X_{1,2}),203(X_{1,2}),205(X_{1,2}),230(N^{*}_{1,2})\end{matrix}$
@@
$D_{4h}$ $B_{1g}$ $k_{x}^{2}-k_{y}^{2}$ $\begin{matrix}126(A_{1,2,3,4}),131(Z_{1,2,3,4},A_{1,2,3,4}),133(Z_{1,2,3,4}),135(Z_{1,2,3,4}),\\ 137(Z_{1,2,3,4}),141(M_{1,2,3,4}),142(M_{1,2,3,4}),223(X^{*}_{1,2,3,4})\end{matrix}$
$D_{4h}$ $B_{2g}$ $k_{x}k_{y}$ $\begin{matrix}125(M_{1,2,3,4},A_{1,2,3,4}),126(M_{1,2,3,4}),129(M_{1,2,3,4},A_{1,2,3,4}),130(M_{1,2,3,4}),\\ 132(Z_{1,2,3,4},A_{1,2,3,4}),133(M_{1,2,3,4}),134(Z_{1,2,3,4},M_{1,2,3,4}),136(Z_{1,2,3,4},A_{1,2,3,4}),\\ 137(M_{1,2,3,4}),138(Z_{1,2,3,4},M_{1,2,3,4}),222(M_{1,2,3,4}),224(M_{1,2,3,4},X^{*}_{1,2,3,4}),227(X^{*}_{1,2,3,4}),228(X^{*}_{1,2,3,4})\end{matrix}$
$D_{4h}$ $A_{2g}$ $k_{x}k_{y}(k_{x}^{2}-k_{y}^{2})$ $124(Z_{1,2},A_{1,2}),126(Z_{1,2}),128(Z_{1,2},A_{1,2}),130(Z_{1,2}),133(A_{1,2}),137(A_{1,2}),222(X^{*}_{1,2})$
@@
$D_{2h}$ $A_{u}$ xyzxyz $\begin{matrix}48(S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),49(R_{1,2},T_{1,2},U_{1,2},Z_{1,2}),\\ 50(R_{1,2},S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),52(U_{1,2},X_{1,2},Z_{1,2}),53(S_{1,2},X_{1,2}),54(T_{1,2},Z_{1,2}),\\ 56(Z_{1,2}),58(Z_{1,2}),60(Y_{1,2}),66(T_{1,2},Z_{1,2}),68(T_{1,2},Z_{1,2}),70(T_{1,2},Y_{1,2},Z_{1,2}),124(R_{1,2}),\\ 125(R_{1,2},X_{1,2}),126(R_{1,2},X_{1,2}),132(R_{1,2}),133(R_{1,2},X_{1,2}),134(R_{1,2},X_{1,2}),\\ 141(X^{*}_{1,2}),142(X^{*}_{1,2}),192(L_{1,2}),201(M_{1,2},X_{1,2}),203(X_{1,2}),230(N^{*}_{1,2})\end{matrix}$
$D_{4h}$ $B_{1u}$ xyzxyz $\begin{matrix}125(A_{3,4},M_{3,4}),126(M_{3,4}),132(A_{1,2},Z_{1,2}),133(M_{1,2}),134(M_{1,2},Z_{1,2}),\\ 136(Z_{1,2}),138(Z_{1,2}),222(M_{3,4}),224(M_{1,2},X^{*}_{1,2}),227(X^{*}_{1,2}),228(X^{*}_{1,2})\end{matrix}$
@@ Appendix F Electronic structure of FeSe
Density-functional theory calculations for bulk FeSe with the $(\pi,\pi,0)$ coplanar magnetic order are carried out using the Full-potential Linearized Augmented Plane Wave (FLAPW) method Weinert et al. (2009), with the scalar-relativistic approximation Koelling and Harmon (1977); MacDonald et al. (1980). The Perdew-Burke-Ernzerhof form of exchange correlation functional Perdew et al. (1996) is used, along with wave function and potential energy cutoffs of 16 and 200 Ry, respectively. A muffin-tin sphere radius of 1.1 Å is applied to both Fe and Se atoms. The superlattice setup is $\sqrt{2}\times\sqrt{2}\times 1$, as indicated by the dotted line in the left panel of Figure 2 of the main text, with experimental lattice parameters ($a=3.765$ Å; $c=5.518$ Å; $z_{\textrm{Se}}=0.26$) Villars (1997). The Brillouin zone is sampled using (15,15,10) and (70,70,50) k-point meshes for self-consistent field calculation and Fermi surface calculation, respectively. The resulting band structure agrees with a previous calculation employing the generalized Bloch theorem for spin spirals Shishidou et al. (2018).
@@ Appendix G Electronic structure of CeNiAsO
For the construction of the minimal low-energy models relevant for CeNiAsO, we perform ab initio calculations within the full-potential local-orbital (FPLO) code Koepernik and Eschrig (1999) (version 22.00-62) in the non-relativistic setting and the LDSA approximation. We use $a=b=4.014537\mathrm{\AA }$, $c=7.90723\mathrm{\AA }$ in space group 129, Ce on 2c with internal coordinates $(1/4,1/4,0.154739)$, Ni on 2b with $(1/4,1/4,1/2)$, As on 2c with (1/4,1/4,0.351368)(1/4,1/4,-0.351368) and O on 2a with $(1/4,3/4,0)$. The resulting band structure is shown in Fig. 5 which exhibits two pairs of not hybridized bands as checked by calculating the irreps of the corresponding bands.